How much heat is needed to raise the temperature of 50.0 g of water from 4.5°C to 83.0°C

1
Expert's answer
2021-11-16T19:19:40-0500

The heat is


"Q=cm(t_2-t_1)=4200·0.05(83-4.5)=\\=16485\text{ J}/\text{kg}·\text K."